# Deep Learning for Continuous-Time Stochastic Control with Jumps

### v2

@@ D.1 Constant coefficients
We then address a $50$-dimensional optimal consumption-investment problem with jumps, where only GPI-CBU is implemented as GPI-PINN becomes numerically infeasible. Figures 12 and 13 again confirm the accuracy of GPI-CBU for both the value function and the optimal consumption rate $c^{*}(t)$, with the standard deviation across 10 independent runs of GPI-CBU being virtually imperceptible. The optimal wealth allocations $\pi^{*}$, being constant, are not depicted. However, the ${\rm MAE}_{\alpha}$ heatmap in Figure 14 corroborates our method’s accuracy in determining both $c^{*}$ and $\pi^{*}$. We also observe in this Figure 14 that the higher absolute errors tend to occur at the boundaries of the domain.
@@
Figure 12: Value function $V(t,y)$ at $t=0$ for $y\in[0,150]$ (left) and at $y=50$ for $t\in[0,1]$ (right), with $n=50$ stocks. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Blue line: numerical solution from the Runge-Kutta scheme applied to (40)-(41).
@@
Figure 13: Optimal consumption rate $c^{*}(t)$ for $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Blue line: numerical solution from the Runge-Kutta scheme applied to (40)-(41).
@@
Figure 14: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $y\in[0,150]$, obtained from GPI-CBU for the optimal consumption-investment problem with constant coefficents.
@@ D.2 Stochastic coefficients
We then study a more complex optimal consumption-investment problem in a realistic market of the form (34)–(35) with a stochastic interest rate
$dr_{t}=a_{r}(b_{t}-r_{t})dt+\nu_{r}\sqrt{r_{t}}\,\Sigma_{r}^{\top}dW_{t},$
@@
$\frac{dS^{i}_{t}}{S^{i}_{t-}}$ $=\mu^{i}dt+\sqrt{\sigma^{i}_{t}}\,\Sigma^{\top}_{S,i}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\,$,
$d\sigma^{i}_{t}$ $=a^{i}_{\sigma}(b^{i}_{\sigma}-\sigma^{i}_{t})+\nu^{i}_{\sigma}\sqrt{\sigma^{i}_{t}}\,\Sigma^{\top}_{\sigma,i}\,dW_{t},$
@@
$d\lambda^{i}_{t}=a^{i}_{\lambda}(b^{i}_{\lambda}-\lambda^{i}_{t})+\nu^{i}_{\lambda}\sqrt{\lambda^{i}_{t}}\,\Sigma^{\top}_{\lambda,i}\,dW_{t}\,.$
Note that $\Sigma=(\Sigma_{S,1},\ldots,\Sigma_{S,n},\Sigma_{\sigma,1},\ldots,\Sigma_{\sigma,n},\Sigma_{\lambda,1},\ldots,\Sigma_{\lambda,n},\Sigma_{r})^{\top}\in\mathbb{R}^{(3n+1)\times(3n+1)}$ is the (upper) Choleski decomposition of the correlation matrix $\Sigma\Sigma^{\top}$ associated with the $(3n+1)$-dimensional Brownian motions $W$. In this case, the wealth process can still be described as follows
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sqrt{\sigma^{i}_{t}}\ \Sigma_{S,i}^{\top}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right],$
and the strategies should be of the form $c(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $\pi^{i}(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$, with $\alpha_{t}=(\pi^{1}_{t},\ldots,\pi^{n}_{t},c_{t})^{\top}\hskip-2.84526pt\in A:=(0,1)^{n+1}$. Hence, denoting the $(2n+2)$-dimensional process $X^{\alpha}_{t}=(\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$, its dynamics are given by
@@
where
$\mu_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}a_{\sigma}(b_{\sigma}-\sigma_{t})\\ a_{\lambda}(b_{\lambda}-\lambda_{t})\\ a_{r}(b_{r}-r_{t})\\ Y^{\alpha}_{t-}(r_{t}+(\mu-r_{t})^{\top}\pi_{t}-c_{t})\end{array}\right),\,\Sigma_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}\nu_{\sigma}\,\sqrt{\sigma_{t}}\,\Sigma^{\top}_{\sigma}\\ \nu_{\lambda}\,\sqrt{\lambda_{t}}\,\Sigma^{\top}_{\lambda}\\ \nu_{r}\,\sqrt{r_{t}}\,\Sigma^{\top}_{r}\\ Y^{\alpha}_{t-}\sum_{i=1}^{n}\pi^{i}_{t}\sqrt{\sigma^{i}_{t}}\,\Sigma_{S,i}^{\top}\end{array}\right),\,$
@@
$\gamma_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}\bm{0}_{n}\\ \bm{0}_{n}\\ 0\\ Y^{\alpha}_{t-}\pi_{t}^{\top}\mathbf{1}_{n}\end{array}\right),$
with $\mu_{X}(\cdot)\in\mathbb{R}^{2n+2},\Sigma_{X}(\cdot)\in\mathbb{R}^{(2n+2)\times(3n+1)}_{+}$ and $\gamma_{X}(\cdot)\in\mathbb{R}^{2n+2}$. Therefore, the value function
$V(t,x)=\sup_{\alpha}\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\gamma}}{\gamma}ds+\frac{\left(Y^{\alpha}_{T}\right)^{\gamma}}{\gamma}\,\Big|\,X^{\alpha}_{t}=x\right],$
satisfies the following HJB equation for all $(t,x)\in[0,T)\times\mathbb{R}^{n}_{+}\times\mathbb{R}^{n}_{+}\times\mathbb{R}\times\mathbb{R}^{+}$,
@@
$0=\hskip 2.13394pt$ $\partial_{t}V(t,x)-\rho V(t,x)+\sup_{(c,\pi)\in A}\bigg\{\mu_{X}^{\top}\hskip-0.56905pt(t,x)\,\nabla_{x}V(t,x)+\frac{1}{2}\text{Tr}\left[\Sigma_{X}(t,x)\Sigma_{X}^{\top}(t,x)\,\nabla^{2}_{x}V(t,x)\right]$
@@
$\hskip 18.77882pt+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[V(t,\sigma,\lambda,r,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,x)\right]+\frac{(cy)^{\gamma}}{\gamma}\bigg\}\,$, (42)
@@
with terminal condition $V(T,x)=y^{\gamma}/\gamma$. We consider a portfolio of $n=25$ stocks, resulting in a $52$-dimensional value function and a $26$-dimensional control process, with the same parameters as in Section D.1. This version of the consumption-investment problem is significantly more complex than the one discussed in Section D.1, with the value function’s dimensionality increasing from 2 to 52. Consequently, the RK45 numerical method can no longer serve as a reference point to solve the associated HJB equation (42), making it impossible to compute the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics. However, despite this lack of a reference for comparison, GPI-CBU still produces results that appear reasonable. We indeed first observe in Figure 4 (main manuscript) that the losses $\widehat{\mathscr{L}}^{(k)}_{1}$ (21) and $\widehat{\mathscr{L}}^{(k)}_{2}$ (23) converge as the number of epoch $k$ increases. Note that the orange curve in the left plot represents the interior loss (first right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, while the blue curve is the boundary loss (second right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, see Eq. (21).
@@
The following Figures 15 and 16 confirm that the results from GPI-CBU for both the value function and the optimal control are reasonable, as they closely resemble those from Figures 12 and 13. For the value functions of Figure 15, the standard deviations across 10 independent runs of GPI-CBU remain very low, confirming the stability of the approximations. For the optimal consumption rate in Figure 16 (left plot), the standard deviation across the 10 runs tends to be higher for values of time $t$ close to 0, although being still reasonable. Finally, varying the first dimension of the intensity $\lambda^{1}$ mainly impacts the corresponding fraction of wealth $\pi^{1,*}_{t},$ while its effect on the other proportions and consumption rate is more moderate (since arising from the correlation matrix $\Sigma\Sigma^{T}$ of the Brownian motion $W$).
@@
Figure 15: Value function $V(t,x)$ at $t=0$ for $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,y)$ with $y\in[0,150]$ (left) and at $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$ for $t\in[0,1]$ (right). Blue line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$).
@@
Figure 16: Left plot: optimal consumption rate $c^{*}(t,x)$ at $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$ for $t\in[0,1]$ with numerical results from GPI-CBU in blue line and $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Right plot: optimal consumption rate $c^{*}(t,x)$ and first five optimal fractions of wealth $\pi^{*}(t,x)$ at $t=0$ and $x=(0.15\cdot\mathbf{1}_{10},\lambda^{1},\mathbf{1}_{9},0.02,50)$ for $\lambda^{1}\in[0,5]$.

### v3

@@ D.1 Constant coefficients
Next, we address a higher-dimensional optimal consumption-investment problem with jumps ($n=50$ stocks), where only GPI-CBU is implemented since GPI-PINN becomes then numerically infeasible. Figures 12 and 13 again confirm the accuracy of GPI-CBU for both the value function and the optimal consumption rate $c^{*}(t)$, with the standard deviation across 10 independent runs of GPI-CBU being virtually imperceptible. The optimal wealth allocations $\pi^{*}$, being constant, are not depicted. However, the ${\rm MAE}_{\alpha}$ heatmap in Figure 14 corroborates our method’s accuracy in determining both $c^{*}$ and $\pi^{*}$. We also observe that the higher absolute errors tend to occur at the boundaries of the domain.
@@
Figure 12: Value function $V(t,y)$ for $t=0$, $y\in[0,150]$ (left) and $t\in[0,1]$, $y=50$ (right) for the optimal consumption-investment problem (20) with constant coefficients and $n=50$ stocks. Orange dotted lines: results of GPI-CBU with $\pm$ 1 standard deviation given by orange shaded area ($k_{*}=1000$). Blue lines: reference solution from Runge–Kutta applied to (32)-(33).
@@
Figure 13: Optimal relative consumption rate $c^{*}(t)$ for $t\in[0,1]$ for the optimal consumption-investment problem (20) with constant coefficients and $n=50$ stocks. Orange dotted line: results of GPI-CBU with $\pm$ 1 standard deviation given by orange shaded area ($k_{*}=1000$). Blue line: reference solution from Runge–Kutta applied to (32)-(33).
@@
Figure 14: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and $y\in[0,150]$, for GPI-CBU applied to the optimal consumption-investment problem (20) with constant coefficients and $n=50$ stocks.
@@ D.2 Stochastic coefficients
Now, we study a more complex optimal consumption-investment problem in a realistic market of the form (27)–(28) with a stochastic interest rate
$dr_{t}=a_{r}(b_{t}-r_{t})dt+\nu_{r}\sqrt{r_{t}}\,dW^{r}_{t},$
@@
$\frac{dS^{i}_{t}}{S^{i}_{t-}}$ $=\mu^{i}dt+\sigma^{i}_{t}\,dW^{S,i}_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\,$,
for $\sigma^{i}_{t}=\sqrt{v^{i}_{t}}$, where $v^{i}_{t}$ follows
$dv^{i}_{t}$ $=a^{i}_{v}(b^{i}_{v}-v^{i}_{t})+\nu^{i}_{v}\sqrt{v^{i}_{t}}\,dW^{v,i}_{t},$ (34)
@@
$d\lambda^{i}_{t}=a^{i}_{\lambda}(b^{i}_{\lambda}-\lambda^{i}_{t})+\nu^{i}_{\lambda}\sqrt{\lambda^{i}_{t}}\,dW^{\lambda,i}_{t}\,.$
We denote $W=(W^{S,1},\ldots,W^{S,n},W^{v,1},\ldots,W^{v,n},W^{\lambda,1},\ldots,W^{\lambda,n},W^{r})^{\top}$ a $(3n+1)$-dimensional Brownian motion with correlation matrix $\Sigma$. In this case, the wealth process can still be described as follows
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sqrt{v^{i}_{t}}\ dW^{S,i}_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right],$
and the strategies should be of the form $c(t,v_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $\pi^{i}(t,v_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$, with $\alpha_{t}=(c_{t},\pi^{1}_{t},\ldots,\pi^{n}_{t})^{\top}\hskip-2.84526pt\in A:=\mathbb{R}_{+}^{n+1}$. Hence, the dynamics of the $(2n+2)$-dimensional process $X^{\alpha}_{t}=(v_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ are given by
@@
with $\mu_{X}(\cdot)\in\mathbb{R}^{2n+2},\Sigma_{X}(\cdot)\in\mathbb{R}^{(2n+2)\times(3n+1)}_{+}$ and $\gamma_{X}(\cdot)\in\mathbb{R}^{2n+2}$ such that
$\mu_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}a_{v}(b_{v}-v_{t})\\ a_{\lambda}(b_{\lambda}-\lambda_{t})\\ a_{r}(b_{r}-r_{t})\\ Y^{\alpha}_{t-}(r_{t}+(\mu-r_{t})^{\top}\pi_{t}-c_{t})\end{array}\right),\,\ \ \gamma_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}\bm{0}_{n}\\ \bm{0}_{n}\\ 0\\ Y^{\alpha}_{t-}\pi_{t}^{\top}\mathbf{1}_{n}\end{array}\right),$
@@
$\Sigma_{X}(t,X^{\alpha}_{t})=\begin{pmatrix}\mathbf{0}_{n\times n}&D_{v}&\mathbf{0}_{n\times n}&\mathbf{0}_{n\times 1}\\[4.0pt] \mathbf{0}_{n\times n}&\mathbf{0}_{n\times n}&D_{\lambda}&\mathbf{0}_{n\times 1}\\[4.0pt] \mathbf{0}_{1\times n}&\mathbf{0}_{1\times n}&\mathbf{0}_{1\times n}&\nu_{r}\sqrt{r_{t}}\\[4.0pt] D_{Y}&\mathbf{0}_{1\times n}&\mathbf{0}_{1\times n}&0\end{pmatrix}$,
where $D_{v}=\text{diag}\left(\nu_{v}^{1}\sqrt{v^{1}_{t}},\dots,\nu_{v}^{n}\sqrt{v^{n}_{t}}\right)_{n\times n},$ $D_{\lambda}=\text{diag}\left(\nu_{\lambda}^{1}\sqrt{\lambda^{1}_{t}},\dots,\nu_{\lambda}^{n}\sqrt{\lambda^{n}_{t}}\right)_{n\times n}$, $D_{Y}=\left(\pi_{t}^{1}Y^{\alpha}_{t-}\sqrt{v^{1}_{t}},\dots,\pi_{t}^{n}Y^{\alpha}_{t-}\sqrt{v^{n}_{t}}\right)_{1\times n}$. Therefore, the value function
$V(t,x)=\sup_{\alpha}\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\delta}}{\delta}ds+e^{-\rho(T-t)}\frac{\left(Y^{\alpha}_{T}\right)^{\delta}}{\delta}\,\Big|\,X^{\alpha}_{t}=x\right],$
satisfies the following HJB equation for all $(t,x)\in[0,T)\times\mathbb{R}^{n}_{+}\times\mathbb{R}^{n}_{+}\times\mathbb{R}\times\mathbb{R}_{+}$,
@@
$0=\hskip 2.13394pt$ $\partial_{t}V(t,x)-\rho V(t,x)+\sup_{(c,\pi)\in A}\bigg\{\mu_{X}^{\top}\hskip-0.56905pt(t,x)\,\nabla_{x}V(t,x)+\frac{1}{2}\text{Tr}\left[\Sigma_{X}(t,x)\Sigma\Sigma_{X}^{\top}(t,x)\,\nabla^{2}_{x}V(t,x)\right]$
@@
$\hskip 18.77882pt+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[V(t,v,\lambda,r,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,x)\right]+\frac{(cy)^{\delta}}{\delta}\bigg\}\,$, (35)
@@
with terminal condition $V(T,x)=y^{\delta}/\delta$. We consider a portfolio of $n=25$ stocks, resulting in a $52$-dimensional value function and a $26$-dimensional control process, with the same parameters as in Section D.1. This version of the consumption-investment problem is significantly more complex than the one discussed in Section D.1, with the value function’s dimensionality increasing from 2 to 52. Consequently, Runge–Kutta can no longer be used to solve the associated HJB equation (35), making it impossible to compute the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics. However, despite this lack of a reference for comparison, GPI-CBU still produces results that appear reasonable. We indeed first observe in Figure 4 (main manuscript) that the losses $\widehat{\mathscr{L}}^{(k)}_{1}$ (14) and $\widehat{\mathscr{L}}^{(k)}_{2}$ (16) converge as the number of epoch $k$ increases. Note that the orange curve in the left plot represents the interior loss (first right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, while the blue curve is the boundary loss (second right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, see Eq. (14).
@@
The following Figures 15 and 16 confirm that the results of GPI-CBU for both the value function and optimal control are reasonable, as they closely resemble those of Figures 12 and 13. For the value functions of Figure 15, the standard deviations across 10 independent runs of GPI-CBU remain very low, confirming the stability of the approximations. For the optimal consumption rate in Figure 16 (left plot), the standard deviation across the 10 runs tends to be higher for values of time $t$ close to 0, although being still reasonable. Finally, varying the first dimension of the intensity $\lambda^{1}$ mainly impacts the corresponding fraction of wealth $\pi^{*,1}_{t}$, while its effect on the other proportions and consumption rate is more moderate (since arising from the correlation matrix $\Sigma$ of the Brownian motion $W$).
@@
Figure 15: Value function $V(t,x)$ for $t=0$ and $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,y)$ with $y\in[0,150]$ (left) and for $t\in[0,1]$ and $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$ (right), for the optimal consumption-investment problem (20) with stochastic coefficients and $n=25$ stocks. Blue lines: results of GPI-CBU with $\pm$ 1 standard deviation given by orange shaded area ($k_{*}=1000$).
@@
Figure 16: Optimal policy for the optimal consumption-investment problem (20) with stochastic coefficients and $n=25$ stocks. Left plot: optimal consumption rate $c^{*}(t,x)$ for $t\in[0,1]$ and $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$, results of GPI-CBU in blue and $\pm$ 1 standard deviation given by orange shaded area ($k_{*}=1000$). Right plot: optimal consumption rate $c^{*}(t,x)$ and first five optimal fractions of wealth $\pi^{*}(t,x)$ for $t=0$ and $x=(0.15\cdot\mathbf{1}_{10},\lambda^{1},\mathbf{1}_{9},0.02,50)$ for $\lambda^{1}\in[0,5]$.